# From Lengthy to Lucid: A Systematic Literature Review on NLP Techniques for Taming Long Sentences

### v1

@@ 3.1.4. Deep Learning
With the advent of deep learning (DL), several neural extractive and abstractive methods have been introduced. Since sentence compression can be easily formulated as a sequence-to-sequence (seq2seq) tasks, several neural models such as Recurrent Neural Networks (RNNs) (Yu et al., 2018; Chopra et al., 2016; Wenfeng et al., 2017; Zhou et al., 2017; Guan, 2021), Convolutional Neural Networks (CNNs) (Hou et al., 2020), Graph Neural Networks (GNNs) and Transformers (Duan et al., 2019b; West et al., 2019; Guan et al., 2021a) have been proposed. The first neural abstractive approach was introduced in (Rush et al., 2015), where an attention-based contextual encoder combined with a feed-forward neural language model is proposed. However, a feed-forward model might not be capable of capturing all the long-term dependencies of sequential information in an input sentence compared to RNNs.
@@
RNNs are among the most common and widely used neural architectures for sentence compression. More specifically, Chopra et al. (2016) extend the previous feed-forward neural model (Rush et al., 2015) with attention-based RNNs, where the decoder is fed with an additional attention-based input during the text generation. Similarly, Yu et al. (2018) propose an abstractive operation network inspired by the well known pointer-generator network (See et al., 2017). This technique models sentence compression as an editing task where redundant words are removed from the input sentence and new words are created either by rephrasing or copying existing ones. Zhou et al. (2017) use RNNs with a selective gate network that filters the representations based on the first-level representation of the encoder before passing the information to the decoder. Choi et al. (2019) employ a pointer generator network with both event and global attention. The event attention allows to focus on tokens that describe some event, which is especially important in the news domain, while the global attention captures the general meaning of the sentence. Then, event attention and global attention are summed together as a weighted average.  Miao and Blunsom (2016) propose a deep generative model with a variational auto-encoder for both abstractive and extractive settings. First, a latent compression is extracted through the generative model, and the variational auto-encoder is used to generate the final compressed sentence according to this compression.
@@
Other RNNs approaches exploit additional features such as contextual embeddings (Zhou and Rush, 2019), knowledge graphs (Guan, 2021) or features from other datasets. For example, Klerke et al. (2016) use a three-layer bidirectional LSTM and incorporate an eye-tracking corpus to improve compressions with a multi-task objective. Kuvshinova and Khritankov (2019) extend the model presented in (Klerke et al., 2016) by modifying its loss function with the probability score of a generative pre-trained language model. Guan (2021) performs abstractive sentence compression with knowledge graphs and LSTMs with a multi-copy attention mechanism. In addition, Zhou and Rush (2019) propose an unsupervised technique that employs ELMo contextual embeddings with a summarization language model to generate both abstractive and extractive compressions.
@@
Another category of RNN techniques consists of unsupervised techniques. More specifically, Malireddy et al. (2020) introduce an unsupervised neural compression method that uses an encoder-decoder bi-LSTM architecture with a compressor and a reconstructor. The compressor decides which tokens from the input should be dropped based on a linkage loss that penalizes the model if the reconstructor fails to re-generate the sentence. Baziotis et al. (2019) use an unsupervised approach with a compressor and a reconstructor similar to (Malireddy et al., 2020), but instead of optimizing the network for dropping tokens from the input, the reconstructor attempts to rephrase the compressed words. Similarly, Févry and Phang (2018) employ a denoising auto-encoder RNN to reconstruct the input text.
@@
Even though DL techniques are inherently capable of handling more complex tasks such as abstractive generation, they have been also used for developing extractive sentence compressors. For example, Nguyen et al. (2020) enhance a bi-LSTM model with contextual embeddings for deletion-based compression. Soares et al. (2020) demonstrate that employing syntactic features such as POS tags and dependencies with LSTMs can lead to comparable performance across different LSTM architectures, even when using an order of magnitude less data.
@@
Other recurrent architectures involve the use of Gated Recurrent Units (GRUs). More specifically, Mallinson et al. (2018) approach compression using neural machine translation by translating the input sentence into another language and then reverse-translating the output back into the original language. GRUs are also employed in (Nayeem et al., 2019) where an abstractive bi-GRU encoder and a simple GRU decoder are combined with a copying attention mechanism. At the same time, to avoid overly compressed sentences, multiple compressions with different compression rates are generated during inference using a diverse beam search decoding strategy.
@@
CNNs initialized with BERT embeddings were employed for extractive sentence compression in (Hou et al., 2020). Even though the experimental results demonstrate that RNNs have stronger performance, this study shows that CNN-based methods achieve a much faster inference time compared to RNN-based ones.
@@
GNNs have also been proposed for the task of sentence compression, either in an abstractive (Guan et al., 2021b) or extractive setting (Park et al., 2021). More specifically, Park et al. (2021) represent each word’s dependency tree with GNNs to obtain a richer structural vector representation for each word, while Guan et al. (2021b) exploit relations and structural information with GNNs as extracted by FrameNet (Fillmore et al., 2001). FrameNet is a semantic database that contains approximately 300 lexical items and their semantic relations, as well as their underlying semantic structural dependencies.
@@
Transformers have emerged as a powerful model architecture for many NLP tasks, including sentence compression. The performance of Transformer-based architectures in various text generation tasks surpasses that of traditional RNN architectures. Recently, West et al. (2019) apply the principle of Information Bottleneck to both abstractive and extractive sentence compression. The Information Bottleneck method refers to the best trade-off between compression and information loss when compressing a variable $X$ according to relevant information $Y$. In the case of sentence summarization, the variable $X$ represents the generated compression, while the relevant information $Y$ indicates the next sentence of $X$. A pre-trained language model is used to generate the optimal compression for predicting the next sentence. The generated predictions are then used for training a self-supervised abstractive sentence compression model.
@@
Supervised approaches include the original Transformers architecture (Vaswani et al., 2017) as well as BERT-based models. For example, Duan et al. (2019b) employ a contrastive attention mechanism that considers less relevant parts of the input along with the traditional attention mechanism of Transformers architecture. A softmin function ensures that irrelevant parts of the input will be discouraged. Guan et al. (2021a) use a BERT-based model for abstractive sentence compression that exploits the external knowledge from the FrameNet semantic database. B-BOC (Shichel et al., 2021) is an optimization technique that can be used with any sentence compression model (both supervised and unsupervised) to extract the best candidate sentences for compression. This technique can be used for abstractive sentence compression as well.
@@
Transformers are also used within the framework of knowledge distillation for transferring knowledge from a larger model, which is also known as the teacher, to a smaller model, known as the student. More specifically, Sclar et al. (2022) employ a large pre-trained language model (GPT-3) to distill its knowledge to a GPT-2 model through an iterative distillation process without supervision, where in each round the student acts as a teacher. The generated examples are filtered based on different aspects, including fidelity (whether the compression can be inferred from the input), compression rate, and contextual filter (whether the next sentence could be predicted from the output). The smaller model is then fine-tuned on the final compression corpus as generated by the teacher model.
3.1.5. Reinforcement Learning
@@
Reinforcement learning (RL) has been studied extensively within the wider field of text summarization. It is based on the interaction of the agent with the environment through rewards. The objective of the agent is to learn and apply an action policy (in this setting, e.g., deleting tokens) that maximizes the sum of rewards received over time, either from the environment (unsupervised) or from a labeled dataset (supervised). Since RL does not require a set of labeled data, it is typically met in unsupervised (Wang et al., 2018; Ghalandari et al., 2022; Hyun et al., 2022) and semi-supervised (Zhao et al., 2018) settings rather than supervised (Zhang et al., 2021) ones.
@@
The first work to use RL in sentence compression, proposed in (Wang et al., 2018), is based on the Deep Q Network (DQN). DQN is a value-based deep reinforcement learning method that uses neural networks to approximate the expected reward of an action within a given state. This work introduces a DQN-based extractive sentence compression method that incorporates hard constraints from previous IPL approaches (Clarke and Lapata, 2008) to be considered in the reward function. These constraints include grammatical, syntactic, and length rules. The agent iteratively deletes tokens from the input sentence until the termination of the process. Similar to (Wang et al., 2018), SCRL (Ghalandari et al., 2022) is another unsupervised extractive approach, starting from a pretrained transformer encoder, which is fine-tuned for sequence classification via reinforcement learning (policy gradient). An interesting aspect is that it is configurable via reward functions (fluency, similarity to source, compression ratio) that can be tailored to specific use cases. Fluency is computed with a masked language model as the average likelihood (logit) of each token in the generated compression.
@@
Hyun et al. (2022) use an RL model formulated as a Markov decision process to solve abstractive sentence compression. The agent’s rewards include the similarity between the source and the target compression as well as the fluency of the output. The fluency of the generation is computed based on the perplexity of a language model (GPT-2). Low perplexity values indicate high fluency in the output, while higher values indicate low fluency. The backbone of this method is a language model that is pre-trained based on prompt-based text with word-level perturbations and length prompts. The perturbations include shuffles and adding or removing words from the input prompt. The model is trained to reconstruct the initial prompt. At the same time, they employ a controllable multi-summary learning mechanism that is capable of generating multiple summaries of different lengths during inference.
@@
Zhao et al. (2018) use a bi-directional RNN as a policy-gradient network to decide which tokens should be retained to maximize a reward function. The output of the policy network is passed to a syntax-based language model, which serves as an evaluator to assess the readability and the grammatical coherence of the compression. The readability score of the evaluator acts as a reward to the policy network in order to update its weights accordingly. Their policy is not initialized randomly but with pseudo-compressions yielded by unsupervised models.
@@
FAR-ASS (Zhang et al., 2021) exploits RL in a supervised manner to improve the factual consistency of the generated compressions. The model is trained with a mixed-objective reward function that combines both factual accuracy and the longest common subsequence between the generated and target compressions.
3.1.6. Summary
@@
Table 1. Overview of sentence compression methods based on type (extractive, abstractive) and training style (supervised, semi-supervised, weakly-supervised, self-supervised, and unsupervised).
Method
Type Training Style
@@
(Knight and Marcu, 2000; Jing, 2000; Knight and Marcu, 2002a; Oguro et al., 2002; Lin, 2003; Vandeghinste and Tjong Kim Sang, 2004; Vandeghinste and Pan, 2004; McDonald, 2006; Unno et al., 2006; Galley and McKeown, 2007; Nomoto, 2007; Clarke and Lapata, 2007; Wu et al., 2007; Yamangil and Nelken, 2008; Clarke and Lapata, 2008; Zajic et al., 2008; Hirao et al., 2009; Galanis and Androutsopoulos, 2010; Wang et al., 2012; Harashima and Kurohashi, 2012; Zhang et al., 2013b; Yoshikawa et al., 2012; Thadani and McKeown, 2013; Zhang et al., 2013a; Filippova and Altun, 2013; Li et al., 2013; Thadani, 2014; Filippova et al., 2015; Klerke et al., 2016; Wenfeng et al., 2017; Wang et al., 2017; Hasegawa et al., 2017; Kamigaito et al., 2018; Rodrigues et al., 2018; Kuvshinova and Khritankov, 2019; Hou et al., 2020; Nóbrega et al., 2020; Nguyen et al., 2020; Park et al., 2021; Zi et al., 2021; Kamigaito and Okumura, 2020; Soares et al., 2020)
(Filippova and Strube, 2008; Xu and Grishman, 2009; Cordeiro et al., 2009; Févry and Phang, 2018; Wang et al., 2008; Schumann et al., 2020; Malireddy et al., 2020; Ghalandari et al., 2022; Liu et al., 2022a)
(Turner and Charniak, 2005)
(Clarke and Lapata, 2006b)
(Zhao et al., 2018)
(Shichel et al., 2021)
(Cohn and Lapata, 2008, 2009; Napoles et al., 2011a; Galanis and Androutsopoulos, 2011; Aziz et al., 2012; Cohn and Lapata, 2013; Rush et al., 2015; Chopra et al., 2016; Zhou et al., 2017; Li et al., 2018; Yu et al., 2018; Mallinson et al., 2018; Duan et al., 2019b; Choi et al., 2019; Nayeem et al., 2019; Duan et al., 2019a; Li et al., 2020b; Zheng et al., 2020; Zhang et al., 2021; Guan, 2021; Guan et al., 2021b, a)
(Li et al., 2020a)
(Liu et al., 2022b)
(Baziotis et al., 2019; Zhou and Rush, 2019; Sclar et al., 2022)
(Miao and Blunsom, 2016)
(West et al., 2019)
3.2. Datasets
Table 2. Datasets for sentence compression. All datasets concern news-related text. Table is ordered by year.
Datasets Type Size Year Evaluation-mainly Task
Ziff-Davis (Knight and Marcu, 2000) Extractive 1,067 2000 Compression
Vandeghinste and Pan (2004) Extractive - 2004 Compression
BNC (Clarke and Lapata, 2008) Extractive 1,433 2008 Compression
Broadcast (Clarke and Lapata, 2008) Extractive 1,370 2008 Compression
Yamangil and Nelken (2008) Abstractive 380K 2008 Compression
Google (Filippova and Altun, 2013) Extractive 250K 2013 Compression
DUC 2004 (Rush et al., 2015) Abstractive 500 2015 Compression, Document Summarization
Gigaword (Napoles et al., 2012; Rush et al., 2015) Abstractive 3.8M 2015 Compression, Document Summarization
MSR (Toutanova et al., 2016) Abstractive 26K 2016 Compression
Newsroom (Grusky et al., 2018; Ghalandari et al., 2022) Extractive 280 2022 Compression, Document Summarization
3.2.1. Extractive Datasets

### v2

@@ 3.1.4 Deep Learning
With the advent of deep learning (DL), several neural extractive and abstractive methods have been introduced. Since sentence compression can be easily formulated as a sequence-to-sequence (seq2seq) tasks, several neural models such as Recurrent Neural Networks (RNNs) Yu2018operation ; Chopra2016 ; Liu2017Attention ; zhou2017selective ; guan2021knowledge , Convolutional Neural Networks (CNNs) hou2020cnn , Graph Neural Networks (GNNs) and Transformers duan2019contrastive ; west2019bottlesum ; Guan2021Frame have been proposed. The first neural abstractive approach was introduced in Rush2015ASummarization , where an attention-based contextual encoder combined with a feed-forward neural language model is proposed. However, a feed-forward model might not be capable of capturing all the long-term dependencies of sequential information in an input sentence compared to RNNs.
@@
RNNs are among the most common and widely used neural architectures for sentence compression. More specifically, Chopra2016 extend the previous feed-forward neural model Rush2015ASummarization with attention-based RNNs, where the decoder is fed with an additional attention-based input during the text generation. Similarly, Yu2018operation propose an abstractive operation network inspired by the well known pointer-generator network See2017GetNetworks . This technique models sentence compression as an editing task where redundant words are removed from the input sentence and new words are created either by rephrasing or copying existing ones. zhou2017selective use RNNs with a selective gate network that filters the representations based on the first-level representation of the encoder before passing the information to the decoder. choi2019event employ a pointer generator network with both event and global attention. The event attention allows to focus on tokens that describe some event, which is especially important in the news domain, while the global attention captures the general meaning of the sentence. Then, event attention and global attention are summed together as a weighted average.  miao2016language propose a deep generative model with a variational auto-encoder for both abstractive and extractive settings. First, a latent compression is extracted through the generative model, and the variational auto-encoder is used to generate the final compressed sentence according to this compression.
@@
Other RNNs approaches exploit additional features such as contextual embeddings zhou2019 , knowledge graphs guan2021knowledge or features from other datasets. For example, klerke2016improving use a three-layer bidirectional LSTM and incorporate an eye-tracking corpus to improve compressions with a multi-task objective. Kuvshinova2019Improving extend the model presented in klerke2016improving by modifying its loss function with the probability score of a generative pre-trained language model. guan2021knowledge performs abstractive sentence compression with knowledge graphs and LSTMs with a multi-copy attention mechanism. In addition, zhou2019 propose an unsupervised technique that employs ELMo contextual embeddings with a summarization language model to generate both abstractive and extractive compressions.
@@
Another category of RNN techniques consists of unsupervised techniques. More specifically, Malireddy2020 introduce an unsupervised neural compression method that uses an encoder-decoder bi-LSTM architecture with a compressor and a reconstructor. The compressor decides which tokens from the input should be dropped based on a linkage loss that penalizes the model if the reconstructor fails to re-generate the sentence. Baziotis2019 use an unsupervised approach with a compressor and a reconstructor similar to Malireddy2020 , but instead of optimizing the network for dropping tokens from the input, the reconstructor attempts to rephrase the compressed words. Similarly, fevry2018unsupervised employ a denoising auto-encoder RNN to reconstruct the input text. Such models Malireddy2020 have also been used in text summarization, where a pretrained compressor is applied to extracted sentences to improve conciseness in multi-document summarization chali2024transfer .
@@
Even though DL techniques are inherently capable of handling more complex tasks such as abstractive generation, they have been also used for developing extractive sentence compressors. For example, nguyen2020contextual enhance a bi-LSTM model with contextual embeddings for deletion-based compression. soares2020sentence demonstrate that employing syntactic features such as POS tags and dependencies with LSTMs can lead to comparable performance across different LSTM architectures, even when using an order of magnitude less data.
@@
Other recurrent architectures involve the use of Gated Recurrent Units (GRUs). More specifically, Mallinson2018 approach compression using neural machine translation by translating the input sentence into another language and then reverse-translating the output back into the original language. GRUs are also employed in nayeem2019diverse where an abstractive bi-GRU encoder and a simple GRU decoder are combined with a copying attention mechanism. At the same time, to avoid overly compressed sentences, multiple compressions with different compression rates are generated during inference using a diverse beam search decoding strategy.
@@
CNNs initialized with BERT embeddings were employed for extractive sentence compression in hou2020cnn . Even though the experimental results demonstrate that RNNs have stronger performance, this study shows that CNN-based methods achieve a much faster inference time compared to RNN-based ones.
@@
GNNs have also been proposed for the task of sentence compression, either in an abstractive guan2021integrating or extractive setting park2021sentence . More specifically, park2021sentence represent each word’s dependency tree with GNNs to obtain a richer structural vector representation for each word, while guan2021integrating exploit relations and structural information with GNNs as extracted by FrameNet fillmore2001building . FrameNet is a semantic database that contains approximately 300 lexical items and their semantic relations, as well as their underlying semantic structural dependencies. zhang2020dense adopt a graph-based approach that captures higher-level semantic relations with GCN layers and generates abstractive compressions using an LSTM decoder. Finally, SWSum guan2025structure extends this line of research by constructing both a semantic scenario graph and a semantic word-relation graph from FrameNet annotations. Then multiple dual graph-based encoder layers are stacked to iteratively exchange information between the two graphs. The resulting fused graph representations are fed into a transformer decoder to generate compressed abstractive summaries.
@@
Transformers have emerged as a powerful model architecture for many NLP tasks, including sentence compression. The performance of Transformer-based architectures in various text generation tasks surpasses that of traditional RNN architectures. Recently, west2019bottlesum apply the principle of Information Bottleneck to both abstractive and extractive sentence compression. The Information Bottleneck method refers to the best trade-off between compression and information loss when compressing a variable $X$ according to relevant information $Y$. In the case of sentence summarization, the variable $X$ represents the generated compression, while the relevant information $Y$ indicates the next sentence of $X$. A pre-trained language model is used to generate the optimal compression for predicting the next sentence. The generated predictions are then used for training a self-supervised abstractive sentence compression model.
@@
Supervised approaches include the original Transformers architecture Vaswani2017 as well as BERT-based models. For example, duan2019contrastive employ a contrastive attention mechanism that considers less relevant parts of the input along with the traditional attention mechanism of Transformers architecture. A softmin function ensures that irrelevant parts of the input will be discouraged. Guan2021Frame use a BERT-based model for abstractive sentence compression that exploits the external knowledge from the FrameNet semantic database. B-BOC Shichel2021 is an optimization technique that can be used with any sentence compression model (both supervised and unsupervised) to extract the best candidate sentences for compression. This technique can be used for abstractive sentence compression as well. guan2023extract propose a modular abstractive summarization framework which consists of three core components: a) knowledge extraction, where the initial sentence is converted into relation triples, b) content selection, where a RoBERTa-based classifier identifies which triples should be included in the summary and c) rewriting, where a BART-large model re-writes the selected triples into a concise, natural language summary.
@@
Transformers are also used within the framework of knowledge distillation for transferring knowledge from a larger model, which is also known as the teacher, to a smaller model, known as the student. More specifically, SclarReferee2022 employ a large pre-trained language model (GPT-3) to distill its knowledge to a GPT-2 model through an iterative distillation process without supervision, where in each round the student acts as a teacher. The generated examples are filtered based on different aspects, including fidelity (whether the compression can be inferred from the input), compression rate, and contextual filter (whether the next sentence could be predicted from the output). The smaller model is then fine-tuned on the final compression corpus as generated by the teacher model. jung2024impossible show that a smaller Transformer can serve as its own teacher by generating multiple candidate summaries and filtering them based on their quality, developing a distilled summarization model without relying on a LLM.
3.1.5 Reinforcement Learning
@@
Reinforcement learning (RL) has been studied extensively within the wider field of text summarization. It is based on the interaction of the agent with the environment through rewards. The objective of the agent is to learn and apply an action policy (in this setting, e.g., deleting tokens) that maximizes the sum of rewards received over time, either from the environment (unsupervised) or from a labeled dataset (supervised). Since RL does not require a set of labeled data, it is typically met in unsupervised Liangguo2018 ; Ghalandari2022EfficientLearning ; hyun2022generating and semi-supervised Zhao2018 settings rather than supervised Zhang2021Fact ones.
@@
The first work to use RL in sentence compression, proposed in Liangguo2018 , is based on the Deep Q Network (DQN). DQN is a value-based deep reinforcement learning method that uses neural networks to approximate the expected cumulative reward of an action, under the assumption of following the optimal policy thereafter. This work introduces a DQN-based extractive sentence compression method that incorporates hard constraints from previous IPL approaches Clarke2008GlobalApproach to be considered in the reward function. These constraints include grammatical, syntactic, and length rules. The agent iteratively deletes tokens from the input sentence until the termination of the process. Similar to Liangguo2018 , SCRL Ghalandari2022EfficientLearning is another unsupervised extractive approach, starting from a pretrained transformer encoder, which is fine-tuned for sequence classification via reinforcement learning (policy gradient). An interesting aspect is that it is configurable via reward functions (fluency, similarity to source, compression ratio) that can be tailored to specific use cases. Fluency is computed with a masked language model as the average likelihood (logit) of each token in the generated compression.
@@
hyun2022generating use an RL model formulated as a Markov decision process to solve abstractive sentence compression. The agent’s rewards include the similarity between the source and the target compression as well as the fluency of the output. The fluency of the generation is computed based on the perplexity of a language model (GPT-2). Low perplexity values indicate high fluency in the output, while higher values indicate low fluency. The backbone of this method is a language model that is pre-trained based on prompt-based text with word-level perturbations and length prompts. The perturbations include shuffles and adding or removing words from the input prompt. The model is trained to reconstruct the initial prompt. At the same time, they employ a controllable multi-summary learning mechanism that is capable of generating multiple summaries of different lengths during inference.
@@
Zhao2018 use a bi-directional RNN as a policy-gradient network to decide which tokens should be retained to maximize a reward function. The output of the policy network is passed to a syntax-based language model, which serves as an evaluator to assess the readability and the grammatical coherence of the compression. The readability score of the evaluator acts as a reward to the policy network in order to update its weights accordingly. Their policy is not initialized randomly but with pseudo-compressions yielded by unsupervised models.
@@
FAR-ASS Zhang2021Fact exploits RL in a supervised manner to improve the factual consistency of the generated compressions. The model is trained with a mixed-objective reward function that combines both factual accuracy and the longest common subsequence between the generated and target compressions.
3.1.6 Summary
@@
3.2 Datasets
Table 1: Datasets for sentence compression. All datasets concern news-related text. Table is ordered by year.
Datasets Type Size Year Evaluation-mainly
Ziff-Davis Knight2000 Extractive 1,067 2000
vandeghinste2004sentence Extractive - 2004
BNC Clarke2008GlobalApproach Extractive 1,433 2008
Broadcast Clarke2008GlobalApproach Extractive 1,370 2008
yamangil2008mining Abstractive 380K 2008
Google Filippova2013OvercomingCompression Extractive 250K 2013
DUC 2004 Rush2015ASummarization Abstractive 500 2015
Gigaword Napoles2012 ; Rush2015ASummarization Abstractive 3.8M 2015
MSR toutanova-etal-2016-dataset Abstractive 26K 2016
Newsroom Grusky2018Newsroom:Strategies ; Ghalandari2022EfficientLearning Extractive 280 2022
Table 2: Overview of sentence compression methods based on type (extractive, abstractive) and training style (supervised, semi-supervised, weakly-supervised, self-supervised, and unsupervised).
Method Type Training Style
@@ 3.2 Datasets
Knight2000 ; Jing2000 ; Knight2002 ; oguro2002evaluation ; lin2003improving ; vandeghinste2004using ; vandeghinste2004sentence ; McDonald2006 ; unno2006trimming ; Galley2007 ; nomoto2007 ; clarke2007modelling ; wu2007 ; yamangil2008mining ; Clarke2008GlobalApproach ; zajic2008 ; hirao2009syntax ; Galanis2010 ; wang2012two ; harashima2012flexible ; zhang2012research ; yoshikawa2012-sentence ; thadani2013-sentence ; zhang2013chinese ; Filippova2013OvercomingCompression ; li2013document ; thadani2014approximation ; Filippova2015 ; klerke2016improving ; Liu2017Attention ; wang2017syntax ; hasegawa2017japanese ; kamigaito2018higher ; frederico2018multilingual ; Kuvshinova2019Improving ; hou2020cnn ; nobrega2020sentence ; nguyen2020contextual ; park2021sentence ; zi2021som ; kamigaito2020syntactically ; soares2020sentence ; tavakoli2025sentence
Filippova2008 ; xu2009parse ; cordeiro2009unsupervised ; fevry2018unsupervised ; wang2008automatic ; Schumann2020DiscreteExtraction ; Malireddy2020 ; Ghalandari2022EfficientLearning ; liu2022learning
turner2005supervised
clarke2006models
Zhao2018
Shichel2021
juseon2024instructcmp
Cohn2008 ; Cohn2009 ; napoles2011paraphrastic ; Galanis2011ACompressor ; aziz2012cross ; cohn2013 ; Rush2015ASummarization ; Chopra2016 ; zhou2017selective ; li2018ensure ; Yu2018operation ; Mallinson2018 ; duan2019contrastive ; choi2019event ; nayeem2019diverse ; duan2019zero ; li2020keywords ; zheng2020controllable ; zhang2020dense ; Zhang2021Fact ; guan2021knowledge ; guan2021integrating ; Guan2021Frame ; guan2023extract ; Lin2023Multimodal ; yuan2024exploring ; guan2025structure ; tollef2025cross
jing2023vision
li2020explicit
Liu2022Controlling
Baziotis2019 ; zhou2019 ; SclarReferee2022
miao2016language
west2019bottlesum
jung2024impossible
3.2.1 Extractive Datasets